OpenStudy (anonymous):

How would I show that the series: 1 + 1/3 -1/2 +1/5 + 1/7 - 1/4 + 1/9 + 1/11 - 1/6 + ... converges to 3/2 ln 2? (Steps if you wouldn't mind would be helpful in my understanding)

OpenStudy (amistre64):

find a rule for the series

OpenStudy (amistre64):

might have to split it up into 3 sequences tho

OpenStudy (anonymous):

Are you sure that about \[ \frac 3 2 \ln 2\]

OpenStudy (amistre64):

\[\sum_{0}^{inf}\frac{1}{2n+1}-\sum_{0}^{inf} \frac{1}{2n+2}\]

OpenStudy (anonymous):

Yes I am confused also eliassaab , because I don't think it converges to 3/2 ln 2

OpenStudy (anonymous):

@GavinxFiasco at what math class level this question was asked?

OpenStudy (anonymous):

@eliassaab it was asked at first year university level.

OpenStudy (anonymous):

One can write the series as \[ \sum_{k=0}^\infty \left( \frac{1}{4 k+1}+\frac{1}{4 k+3}-\frac{1}{2 k+2}\right) = \sum_{k=0}^\infty \left( \frac{8 k+5}{2 (k+1) (4 k+1) (4 k+3)}\right) \] And if you use Wolfram Alpha, you fine that the answer is \[ \frac 32 \ln 2\]

OpenStudy (anonymous):

oh thank you this makes alot of sense